After-Tax Profit Targets Olivian Company wants to earn $540,000 in net (after-tax) income next year. Its product is priced at $350 per unit. Product costs include: Direct materials $105.00 Direct labor $77.00 Variable overhead $17.50 Total fixed factory overhead $420,000 Variable selling expense is $14 per unit; fixed selling and administrative expense totals $270,000. Olivian has a tax rate of 40 percent.

Required: 1. Calculate the before-tax profit needed to achieve an after-tax target of $540,000. $

2. Calculate the number of units that will yield operating income calculated in Requirement 1 above. If required, round your answer to the nearest whole unit. units Hide Feedback Partially Correct Check My Work Feedback 1. After-tax Profit/(1-Tax Rate) = Before-Tax Profit.

2. The required number of units = (Fixed Costs + Target Profit)/Contribution Margin per Unit. Hide

3. Prepare an income statement for Olivian Company for the coming year based on the number of units computed in Requirement 2. Do NOT round interim calculations and, if required, round your answer to the nearest dollar. Olivian Company Income Statement For the Coming Year Total $ $ $ $ Hide Feedback Partially Correct Check My Work Feedback 3. See Cornerstone 16.4. Use a contribution margin format income statement.

4. What if Olivian had a 35 percent tax rate? Would the units sold to reach a $540,000 target net income be higher or lower than the units calculated in Requirement 2? Calculate the number of units needed at the new tax rate. In your calculations, round before-tax income to the nearest dollar. Round your answer to the nearest whole unit. units
Answer #1

Working:

Sales price per unit 350.00
Less variable expenses
Direct materials 105.00
Direct labor 77.00
Variable overhead 17.50
Variable selling expense 14.00
Total variable expenses 213.50
Contribution margin per unit 136.50
Fixed costs:
Fixed factory overhead 420000
Fixed selling and administrative expense 270000
Total fixed costs $ 690000

1. Before-tax profit = After-tax profit/(1 - Tax rate) = $540000/(1 - 0.40) = $540000/0.60 = $900000

2. The required number of units = (Fixed costs + Target profit)/Contribution margin per unit = ($690000 + $900000)/$136.50 = $1590000/$136.50 = 11648.35 = 11648 units

3.

Olivian Company
Income Statement
For the Coming Year
Total
Sales revenue (11648 x $350) 4076800
Less: Variable expenses (11648 x $213.50) 2486848
Contribution margin 1589952
Less: Fixed expenses 690000
Net operating income $ 899952

4. Lower

Before-tax profit = After-tax profit/(1 - Tax rate) = $540000/(1 - 0.35) = $540000/0.65 = $830769.23 = $830769

The required number of units = (Fixed costs + Target profit)/Contribution margin per unit = ($690000 + $830769)/$136.50 = $1520769/$136.50 = 11141.16 = 11141 units